Company XYZ operates in the highly competitive electronics industry. Price for its product is stable at $50. This means that P=MR=$50 in this market. The company manager estimates the relevant total cost function is: TC=78000 +18Q+0.02Q^2

a) Derive the MC function

b) Calculate the output level that will maximize the XYZ profit.

c) Calculate the maximum profit.

d) Does non-price strategy effective in highly competitive industries?
Answer #1

A.

TC=78000 + 18Q + 0.02Q^2

Differentiation of above equation with respect to Q, will give MC.

MC = 18 + .04Q

---------

B

For maximization of profit,

MR = MC

50 = 18 + .04Q

Q = (50-18)/.04

Q = 800 units

----------

C.

Maximum profit = 50*800 - (78000 + 18*800 + 0.02*800^2)

Maximum profit = -$65200

So, there is a loss of $65200.

-------------

D.

Non-price strategy, can be used by any firm such as innovation and creativity, but it will not be effective to raise price in highly competitive industry as firms are price takers. If firms are price makers, such as monopolistic market and oligopoly market, then non-price strategy will be more effective.